Recent Log Entries

LEM South Linac - failure

Log Number: 3813894Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the South Linac on Thu Jul 9 19:52:00 2020. The operation terminated after 0:03:50 with the following results.
Cavity summarySuccess: 197 Abort: 3 Fail: 0

LEM North Linac - failure

Log Number: 3813893Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the North Linac on Thu Jul 9 19:43:03 2020. The operation terminated after 0:03:43 with the following results.
Cavity summarySuccess: 186 Abort: 14 Fail: 0

LEM North Linac - failure

Log Number: 3813890Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the North Linac on Thu Jul 9 19:14:29 2020. The operation terminated after 0:01:27 with the following results.
Cavity summarySuccess: 170 Abort: 30 Fail: 0

LEM North Linac - failure

Log Number: 3813889Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the North Linac on Thu Jul 9 19:11:46 2020. The operation terminated after 0:01:20 with the following results.
Cavity summarySuccess: 134 Abort: 66 Fail: 0

1L07 Cav 4 High Klystron Reflected Power

Log Number: 3813885Logbook(s): Entry Makers: deirBody: 

Cavs 2, 5, 6 in zone 1L07 seem to be dropping due to cav 4. Turning cav 4 on at the GSET of 9.8 causes it to drop after a minute or so. I turned on 2,5,6 and left 4 off, things seemed stable. I reduced GSET in cav 4 to 9 and let it sit. Gradually increased it to 9.2. Shortly after, spikes in R174CRRP,CRRRPP,CRRPQ,CRRPI seem to have caused cavs 2,5,6 to drop to SEL.

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 3813884Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: aiken
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C01<

1L25 Cav 8 Troubles

Log Number: 3813881Logbook(s): Entry Makers: deirBody: 

Can't seem to keep this cavity on. The zone is stable as long as this cavity is off. I tried doing a phase optimization but that didn't help because it wasn't making any GMES in SEL mode. Tried to do a phase sweep (before I realized there was no GMES) and the liquid levels fluctuated and dropped the zone, Fig 1.

Pages